If (x+4),(x+12),(x-1) and (x+5) are proportion then find the value of “x”

Grade:7

1 Answers

Harshit Singh
3 years ago
Dear student

x+4/x+12 = x – 1/x+5
by cross multiplication
(x+4)(x+5) = ( x – 1)(x+12)
x^2 + 9x + 20 = x^2 + 11x –12

2x= 32

x= 16

Thanks